OpenStudy (anonymous):

Okay, I'm attempting to simplify an expression with radicals in it so I can find the limit. I'll type the function below.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sqrt{x+\Delta x} -\sqrt{x}/(\Delta x)\]

OpenStudy (anonymous):

How would you work that?

OpenStudy (campbell_st):

you need to rationalize the numerator multiply by \[\frac{\sqrt{x + \Delta x} + \sqrt{x}}{\sqrt{x + \Delta x} + \sqrt{x}}\]

OpenStudy (campbell_st):

so you end up with \[\lim_{\Delta x \rightarrow 0} \frac{\sqrt{ x + \Delta x} - \sqrt{x}}{\Delta x} \times \frac{\sqrt{ x + \Delta x} + \sqrt{x}}{\sqrt{x + \Delta x} + \sqrt{x}}\] this will result in the difference of 2 squares in the numerator

OpenStudy (anonymous):

Oh shoot, now it's seems obvious. I just needed to find the conjugate of the whole numerator instead of just one part, yes?

OpenStudy (campbell_st):

thats correct... you'll end up with \[\lim_{\Delta x \rightarrow 0} \frac{x + \Delta x - x}{\Delta x(\sqrt{ x + \Delta x} + \sqrt{x}}\] eliminate the common factor and in the denominator as delta x approaches 0 you'll end up with the solution.

OpenStudy (anonymous):

So would I just leave the simplified version ends up being: \[\lim_{x \rightarrow 0} \frac{ 1 }{\sqrt{x + \Delta x} + \sqrt {x} }\]

OpenStudy (campbell_st):

correct... just collect the like terms and delta x disappears

OpenStudy (anonymous):

And then use substitution to replace delta x with 0? (To find the limit?)

OpenStudy (campbell_st):

well substitute 1st.... then you can collect the like terms... \[\frac{1}{\sqrt{x} + \sqrt{x}}\]

OpenStudy (anonymous):

Yea. So the limit ends up being \[\frac{ 1 }{ 2\sqrt{x} }\] Correct?

OpenStudy (campbell_st):

thats it.... thats differentiation by 1st principals.

OpenStudy (anonymous):

AH! Thank you! I'm going to close the question now.

OpenStudy (campbell_st):

hope it helped